For the Equation 4X^2-100=0, did I do my work right? 4X^2-100=0
+100 +100
4X^2=100
/4 /4
X^2=25
My teacher said that if it was a prime number or if it had no perfect square you could leave it as it is is this true and is my answer right? Thank you

Answer by checkley77(12844)

4X^2-100=0
+100 +100
4X^2=100
/4 /4
X^2=25
YOUR ANSWER IS CORRECT SO FAR BUT:
NO YOU CAN'T QUIT NOW BECAUSE 25 IS A PERFECT SQUARE AND THE ANSWER IS:
X=SQRT25
X=5 ANS.